A particle of mass m is attracted to a force center with the force of magnitude k/ r2. Use plane polar coordinates and find Hamilton's equations of motion.


1
Expert's answer
2021-12-06T09:41:47-0500

"F=\frac k{r^2},"

"L=T-U=\frac{m\dot r^2}2-Fr,"

"\frac{\partial L}{\partial r}-\frac d{dt}\frac{\partial L}{\partial \dot r}=0,"

"mr\dot r-\frac k{r^2}\ddot r=0."